# Common Task Framework For a Critical Evaluation of Scientific Machine Learning Algorithms

### v3

@@ 2.1.4 Test 4: Parametric Generalization (2 scores)
Finally, the ability of a model to generalize to different parameter values is evaluated. For this case, the model’s ability to interpolate and extrapolate to new parameter regimes is considered with noise-free data and noisy data respectively. The interpolation and extrapolation are each their own score. This gives a total of four scores that evaluate parametric dependence.
@@ 2.2 Dynamical System: Lorenz
LSTM hochreiter1997lstm 64.54 ($\pm$ 0.00) 99.34 ($\pm$ 0.18) 52.37 ($\pm$ 12.34) 97.42 ($\pm$ 0.10) 50.75 ($\pm$ 28.41) 96.44 ($\pm$ 0.10) 72.13 ($\pm$ 3.54) 66.61 ($\pm$ 6.43) 32.75 ($\pm$ 16.91) 36.39 ($\pm$ 0.00) 29.57 ($\pm$ 14.47) 80.58 ($\pm$ 0.00) 60.11 ($\pm$ 0.00)
@@
Spacetime zhang2023spacetime 38.79 ($\pm$ 0.00) 19.26 ($\pm$ 0.00) 83.52 ($\pm$ 15.01) -100.00 ($\pm$ 0.00) 12.16 ($\pm$ 66.22) 39.32 ($\pm$ 0.00) 56.00 ($\pm$ 8.78) 28.28 ($\pm$ 0.00) 23.68 ($\pm$ 34.66) 33.12 ($\pm$ 0.00) 0.00 ($\pm$ 0.00) 77.32 ($\pm$ 0.00) 54.09 ($\pm$ 0.00)
KAN liu2025kankolmogorovarnoldnetworks 35.32 ($\pm$ 0.00) 82.89 ($\pm$ 26.43) 20.53 ($\pm$ 58.61) 96.19 ($\pm$ 0.06) 55.20 ($\pm$ 3.28) 93.00 ($\pm$ 0.01) -76.80 ($\pm$ 32.81) 50.52 ($\pm$ 8.64) -2.45 ($\pm$ 34.69) 33.68 ($\pm$ 4.41) -31.47 ($\pm$ 38.67) 41.69 ($\pm$ 3.90) 60.85 ($\pm$ 0.17)
SINDy sindy; ensemblesindy 33.08 ($\pm$ 0.00) 81.83 ($\pm$ 0.00) 36.00 ($\pm$ 0.00) 36.85 ($\pm$ 0.00) 69.92 ($\pm$ 2.74) -29.22 ($\pm$ 0.00) -18.27 ($\pm$ 0.00) 55.82 ($\pm$ 8.48) 15.20 ($\pm$ 28.53) 36.83 ($\pm$ 2.90) 14.56 ($\pm$ 23.22) 82.38 ($\pm$ 0.00) 15.07 ($\pm$ 0.00)
OptDMD askham2018variable 20.62 ($\pm$ 0.00) 52.08 ($\pm$ 0.00) -67.33 ($\pm$ 0.00) 55.51 ($\pm$ 0.00) 1.33 ($\pm$ 0.00) 56.85 ($\pm$ 0.00) -64.13 ($\pm$ 0.00) 59.11 ($\pm$ 0.00) 8.67 ($\pm$ 0.00) 42.15 ($\pm$ 0.00) -15.73 ($\pm$ 0.00) 59.23 ($\pm$ 0.00) 59.68 ($\pm$ 0.00)
@@
ODE-LSTM coelho2024odelstm 15.28 ($\pm$ 0.00) 80.09 ($\pm$ 0.27) 0.48 ($\pm$ 50.76) 88.65 ($\pm$ 0.06) -31.46 ($\pm$ 22.33) 52.18 ($\pm$ 0.42) -5.86 ($\pm$ 11.52) 1.71 ($\pm$ 7.42) 49.55 ($\pm$ 7.15) 6.37 ($\pm$ 3.63) 8.52 ($\pm$ 3.08) -54.07 ($\pm$ 0.00) -12.76 ($\pm$ 24.73)
@@ A.1 Dataset Files and Evaluation Metrics
LSTM [hochreiter1997lstm] 78.07 ($\pm$ 7.44) 41.33 ($\pm$ 12.60) 70.34 ($\pm$ 0.00)
@@
Spacetime [zhang2023spacetime] 42.05 ($\pm$ 18.00) 21.27 ($\pm$ 8.66) 65.70 ($\pm$ 0.00)
@@
SINDy [sindy, ensemblesindy] 29.52 ($\pm$ 0.46) 30.60 ($\pm$ 15.78) 48.73 ($\pm$ 0.00)
@@
ODE-LSTM [coelho2024odelstm] 30.68 ($\pm$ 14.23) 16.54 ($\pm$ 5.32) -33.42 ($\pm$ 12.36)
@@
SINDy [sindy, ensemblesindy] -24.80 ($\pm$ 0.00) 16.39 ($\pm$ 0.00) 10.26 ($\pm$ 0.00)

### v4

@@ 2.1.4 Test 4: Parametric Generalization (2 scores)
Finally, the ability of a model to generalize to different parameter values is evaluated. For this case, the model’s ability to interpolate and extrapolate to new parameter regimes is considered with noise-free data. The interpolation and extrapolation are each their own score, resulting in two scores that evaluate parametric dependence.
@@ 2.2 Dynamical System: Lorenz
LSTM hochreiter1997lstm 64.54 ($\pm$ 0.00) 99.34 ($\pm$ 0.18) 52.37 ($\pm$ 12.34) 97.42 ($\pm$ 0.10) 50.75 ($\pm$ 28.41) 96.44 ($\pm$ 0.10) 72.13 ($\pm$ 3.54) 66.61 ($\pm$ 6.43) 32.75 ($\pm$ 16.91) 36.39 ($\pm$ 2.05) 29.57 ($\pm$ 14.47) 80.58 ($\pm$ 0.00) 60.11 ($\pm$ 0.00)
@@
Spacetime zhang2023spacetime 39.46 ($\pm$ 0.00) 19.26 ($\pm$ 0.00) 83.52 ($\pm$ 15.01) 46.79 ($\pm$ 0.01) 12.16 ($\pm$ 66.22) 39.32 ($\pm$ 0.00) 56.00 ($\pm$ 8.78) 28.28 ($\pm$ 0.00) 23.68 ($\pm$ 34.66) 33.12 ($\pm$ 0.00) 0.00 ($\pm$ 0.00) 77.32 ($\pm$ 0.00) 54.09 ($\pm$ 0.00)
OptDMD askham2018variable 20.62 ($\pm$ 0.00) 52.08 ($\pm$ 0.00) -67.33 ($\pm$ 0.00) 55.51 ($\pm$ 0.00) 1.33 ($\pm$ 0.00) 56.85 ($\pm$ 0.00) -64.13 ($\pm$ 0.00) 59.11 ($\pm$ 0.00) 8.67 ($\pm$ 0.00) 42.15 ($\pm$ 0.00) -15.73 ($\pm$ 0.00) 59.68 ($\pm$ 0.00) 59.23 ($\pm$ 0.00)
@@
ODE-LSTM coelho2024odelstm 11.85 ($\pm$ 0.00) 80.09 ($\pm$ 0.27) 0.48 ($\pm$ 50.76) 88.65 ($\pm$ 0.06) -31.46 ($\pm$ 22.33) 52.18 ($\pm$ 0.42) -47.01 ($\pm$ 54.73) 1.71 ($\pm$ 7.42) 49.55 ($\pm$ 7.15) 6.37 ($\pm$ 3.63) 8.52 ($\pm$ 3.08) -54.07 ($\pm$ 0.00) -12.76 ($\pm$ 24.73)
@@ A.1 Dataset Files and Evaluation Metrics
LSTM [hochreiter1997lstm] 78.07 ($\pm$ 7.44) 41.33 ($\pm$ 9.96) 70.34 ($\pm$ 0.00)
@@
Spacetime [zhang2023spacetime] 42.84 ($\pm$ 15.00) 21.27 ($\pm$ 8.66) 65.70 ($\pm$ 0.00)
@@
SINDy [sindy, ensemblesindy] 1.73 ($\pm$ 0.00) 30.60 ($\pm$ 15.78) 48.73 ($\pm$ 0.00)
@@
ODE-LSTM [coelho2024odelstm] 23.82 ($\pm$ 21.43) 16.54 ($\pm$ 5.32) -33.42 ($\pm$ 12.36)
@@
SINDy [sindy, ensemblesindy] -20.35 ($\pm$ 0.00) 16.39 ($\pm$ 0.00) 10.26 ($\pm$ 0.00)